OpenStudy (anonymous):

Is ln(a(1+x)^(-1-a)) = (-1-a)*(ln(a)+ln(1+x)) True ? or is ln(a(1+x)^(-1-a)) = ln(a) + (-1-a)*ln(1+x)) True?

Parth (parthkohli):

Is this the question?\[\ln \left(a \times (1 + x)^{-1-a} \right)\]Or is this the one?\[\ln \left(\left(a (1 + x)\right)^{-1-a}\right)\]

OpenStudy (anonymous):

The first one :)

OpenStudy (anonymous):

when you write it like that it's ln(a(1+x)^(-1-a)) = ln(a) + (-1-a)*ln(1+x))

OpenStudy (anonymous):

i assume :P

Parth (parthkohli):

Right.

OpenStudy (anonymous):

Thanks for your help :)

Parth (parthkohli):

No problem!